The market for gravel has been estimated to have these supply and demand relationships:

Demand: Qd = 10,000 - 100P

Supply: Qs = 100P - 1,000

where P represents price per unit in £, and Q represents sales per week in tons. Determine the amount of shortage or surplus that would develop at P = £40/ton.
1
Expert's answer
2020-11-16T06:59:33-0500

At "P = \pounds 40/ton"

Demand, "Q_{d} = 10,000-100\times40=10,000-4000=6,000"

Supply, "Q_{s} = 100\times40-1000=4000-1000=3000"

Demand is greater than supply

the shortage is "6000-3000=3000"